OpenStudy (mindblast3r):

30-60-90 triangle question

TheSmartOne (thesmartone):

Post away

OpenStudy (mindblast3r):

@thesmartone

OpenStudy (mindblast3r):

@happyrosy

OpenStudy (mindblast3r):

@Nnesha

OpenStudy (anonymous):

This is more properties of a 30-60-90 triangle than Pythagorean theorem.

TheSmartOne (thesmartone):

Do you know the relationship of sides of a 30-60-90 triangle?

OpenStudy (mindblast3r):

YES

TheSmartOne (thesmartone):

Can you draw it using the draw tool?

OpenStudy (mindblast3r):

ok i already know the answer and it showed me it but can you explain how the answer was worked out?

TheSmartOne (thesmartone):



OpenStudy (mindblast3r):

ok, but i don't know how to solve it

TheSmartOne (thesmartone):

They gave you the side opposite the right angle as \(\sf 8\sqrt{3}\) And so if we look at the drawing I posted, and correlate the sides to solve for a we get: \(\sf 2a = 8\sqrt{3}\) \(\sf\Large a=\frac{8\sqrt{3}}{2}=4\sqrt{3}\)

TheSmartOne (thesmartone):

and you can see that the side that they want is the side opposite the angle 60 In the drawing I posted, it has a value of \(\sf a\sqrt{3}\) and we calculated a as \(\sf 4\sqrt{3}\) So if we plug in the value of a, we get: \(\sf 4 \sqrt{3}\times \sqrt{3}=4 \times 3 = 12\)

OpenStudy (anonymous):

The side opposite 30 degrees is always half of the hypotenuse. So if the hypotenuse is \[8\sqrt{3}\] then the side opposite to the 30 degees would be \[\frac{8\sqrt{3}}{2}\] which is \[4\sqrt{3}\] And the 60 degrees side is always the 30 degrees side times sqrt 3. So it would be \[4\sqrt{3}*\sqrt{3}\] which is the same as \[4\sqrt{9}\] which is \[4*3\] which obviously get you 12.